### v1

# Accelerating Speculative Decoding using Dynamic Speculation Length

Abstract
Speculative decoding is a promising method for reducing the inference latency of large language models. The effectiveness of the method depends on the speculation length (SL)—the number of tokens generated by the draft model at each iteration. The vast majority of speculative decoding approaches use the same SL for all iterations. In this work, we show that this practice is suboptimal. We introduce DISCO, a DynamIc SpeCulation length Optimization method that uses a classifier to dynamically adjust the SL at each iteration, while provably preserving the decoding quality. Experiments with four benchmarks demonstrate average speedup gains of 10.3% relative to our best baselines. [^1] Code will be available upon publication.
Accelerating Speculative Decoding using Dynamic Speculation Length
Jonathan Mamou  Oren Pereg  Daniel Korat Moshe BerchanskyNadav TimorMoshe WasserblatRoy Schwartz Intel Labs, Israel Weizmann Institute of Science, Israel School of Computer Science & Engineering, Hebrew University of Jerusalem Correspondence: jonathan.mamou@intel.com
1 Introduction
Large language models (LLMs) generate tokens in an autoregressive manner, incurring slow and costly generation. Speculative decoding methods Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023) mitigate this issue by splitting the inference of a given target model into two steps. In the first step, a faster draft model generates tokens autoregressively. In the second step, the target model validates all the draft tokens simultaneously. The validation procedure accepts or rejects the draft tokens (Fig. 1). This approach accelerates the generation process by producing multiple tokens in a single forward pass of the target model, while maintaining the decoding quality of the target model.
The effectiveness of speculative decoding in expediting the token generation process depends on the speculation length (SL) parameter, which pertains to the number of tokens generated by the draft model before each validation step (Chen et al., 2023). The choice of SL has a large effect on latency: lower static (constant) SL values may incur a lower number of accepted draft tokens per speculative iteration, which leads to more target model validations than necessary. In contrast, higher static SL may result in a large number of ignored draft generations in case of an early draft-target mismatch (Fig. 1).
Figure 1: An illustration of a single speculative decoding iteration with static speculation length ($SL$) $=5$. A draft model generates $SL$ tokens autoregressively. The target model validates them all in parallel, and accepts the first three tokens. In this case, $SL$ is too high (as the last two draft tokens are never used). $SL$ can also be too low (e.g., $SL=2$ here), which would lead to too many calls to the target model.
The vast majority of speculative decoding approaches Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023); Cai et al. (2024); Li et al. (2024) employ a static SL, i.e., a constant number of draft tokens are generated in each speculative iteration. In this work, we show that this approach is suboptimal. We start by introducing a dynamic oracle SL, which sets the SL to be the exact number of draft generations until the first mismatch in each speculative iteration. We observe a high variance in the optimal SL across the speculative iterations. This finding demonstrates the potential of dynamically adapting the speculation length.
We then propose DISCO, a dynamic speculation length optimization method. DISCO uses draft model features to train a lightweight classifier that predicts the compatibility between a generated draft token and the target token. At inference time, the classifier determines whether the draft model should continue generating the next token, or halt and switch to the target model for validation.
We evaluate DISCO’s effectiveness in code generation, text summarization and instruction tasks. Our results demonstrate average speedup gains of 10.3% and 31.4% relatively to the best static and dynamic SL baselines, respectively, while maintaining the same generated text. Our classifier also transfers well across tasks from the same category: training it on one task and using it on another leads to similar gains over our baselines.
Figure 2: The improvement factor (IF) as a function of the speculation length ($\gamma$) for three different acceptance rate ($\alpha$) values assuming fixed cost coefficient ($c=0.05$). The IF value largely depends on both $\alpha$ and $\gamma$.
2 Background: Speculative Decoding
Speculative decoding Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023) expedites LLM generation while ensuring no accuracy loss by dividing it into two stages. In the first stage, a fast but less accurate draft model $M_{D}$ autoregressively generates a sequence of tokens. In the second stage, a large but more accurate target model $M_{T}$ conducts parallelized verification over the generated draft tokens. This process allows the model to potentially produce multiple tokens per target forward pass.
Speculation length (SL)
The effectiveness of speculative decoding in accelerating the token generation process relies heavily on the SL parameter, which determines how many tokens are generated by the draft model before each validation step. The effect of the SL on the overall speedup is subject to a tradeoff; higher SLs potentially reduce the number of target model validations, but also increase the number of redundant draft generations (Fig. 1).
The majority of speculative decoding approaches Leviathan et al. (2023); Miao et al. (2024); Chen et al. (2023); Cai et al. (2024); Li et al. (2024) use a static SL—the same number of draft tokens are generated per speculative iteration.222 Some practical work apply heuristics for dynamic SL adaptation, increasing/decreasing SL when all/some draft tokens are accepted/rejected in the previous iteration (https://huggingface.co/blog/assisted-generation). Chen et al. (2023) explored various static SLs, across different target-draft model pairs and tasks, and empirically showed that as the SL rises, the overall speedup increases until reaching a certain threshold, beyond which it either levels off or even regresses.
To study the effect of different static SL values, Leviathan et al. (2023) defined the improvement factor (IF) as the expected token generation speedup. Formally:
$IF={\frac{1-\alpha^{\gamma+1}}{(1-\alpha)({\gamma}c+1)}}$ (1)
where $\alpha$ denotes the acceptance rate, indicating the expected probability of a draft token to be accepted by the target model; $c$ represents the cost coefficient, indicating the ratio between the walltime of a forward pass run of the draft model $M_{D}$ and the wall time of a forward pass run of the target model $M_{T}$; and $\gamma$ represents the static SL value. [^3] IF computation assumes enough compute resources for increased concurrency as $\gamma$ rises.
Figure 2 shows the IF as a function of $\gamma$ for three different values of $\alpha$ and a fixed $c$ value. While both $\alpha$ and $c$ are important to the selection of the target-draft model pair, finding the optimal $\gamma$ is fundamental to the effectiveness of the system.
3 Dynamic Speculation Length
The IF function is based on the simplifying assumption that the probability of accepting draft tokens by the target model is i.i.d. Nevertheless, in practical scenarios, different tokens may have varying levels of predictability, rendering the simplifying i.i.d. assumption unsound. Therefore, relying on a static SL might not be optimal, as certain speculative iterations may benefit from generating more tokens than others. Below we consider an oracle experiment, which applies the optimal dynamic $\gamma$ value at each iteration. We then propose a method for dynamically setting $\gamma$, showing that it strongly outperforms a static selection method for any choice of a static SL.
Assessing the optimal dynamic $\gamma$ value
We start by employing an oracle for detecting the ‘ground truth’ value of $\gamma$ for each speculative iteration. The oracle uses the draft model to autoregressively generate tokens until a mismatch occurs between the predicted tokens of the draft and target models. This process is repeated for each speculative iteration, ultimately returning the optimal (maximum) number of accepted draft tokens per iteration. The mismatch between the tokens is determined by using the rejection sampling algorithm introduced by Leviathan et al. (2023) with zero temperature. This oracle fulfills the speculative decoding potential: generating the maximal number of valid draft tokens at each iteration, and making a minimal number of calls to both draft and target model.
Figure 3 shows the oracle SL values across the speculative iterations for one MBPP example. Comparing with static SL, we observe a lower number of oracle draft and target forward passes. Figure 4 shows the average oracle SL over the speculative iterations for Alpaca dataset. We observe on both figures a high variance of oracle SL values, implying that setting a predefined static SL is likely to be suboptimal. See App. A for further analysis.
Figure 3: Oracle and static SL values for different speculative iterations on one MBPP example. For static SL, we run 38 target forward passes and 192 draft forward passes, while for oracle SL, we only run 27 target forward passes and 129 draft forward passes. We observe a high variance of oracle SL values.
Figure 4: Bar chart of the average oracle SL over the normalized index of the speculative iterations for the Alpaca dataset. We observe a high variance of oracle SL values.
DynamIc SpeCulation length Optimization
We introduce DISCO, a simple method for dynamically setting the SL value at each iteration. To estimate the correct SL value at each step, we employ a simple SL classifier. Our classifier decides, at each draft generation step, whether the draft model $M_{D}$ should proceed to generate the next token or switch to the target model $M_{T}$ for verification. The classifier takes as inputs the probability vector of the draft model ($y^{D}_{i}$) and the token position ($i$), generating a confidence score ($C_{i}$) used for the decision-making as follows:
$C_{i}=FFN(Concat(Top_{k}(y^{D}_{i}),Ent(y^{D}_{i}),i))$
where $Top_{k}()$ selects the top $k$ values [^4] We use $k=10$ in all experiments. and $Ent()$ is the entropy function.
[…]
4 Experiments
Datasets
We evaluate our method on four datasets spanning three tasks: code generation, using MBPP Austin et al. (2021) and HumanEval Chen et al. (2021); text summarization using CNN-DailyMail Nallapati et al. (2016); and instruction-following using Alpaca Taori et al. (2023). We use the training set for training the SL classifier and the validation set for setting threshold $\tau$ and $SL_{max}$ hyperparameters. For HumanEval, which has no training and validation sets, we evaluate transfer learning from MBPP. See App. B for more details on the datasets and prompts.
Models
To evaluate speculative decoding on code generation tasks, we use the Starcoder model family Li et al. (2023)—15B for target and 168M for draft. For our text generation tasks, we use Vicuna models—13B as target Chiang et al. (2023) and 68M as draft Yang et al. (2024). Throughout the experiments we use the rejection sampling scheme Leviathan et al. (2023) with the temperature set to zero. Exact model names and versions are provided in App. C.
SL classifier training
For minimal overhead, we use a shallow 2-layer FFN classifier. We train the SL classifier to predict the conformity between $M_{D}$ and $M_{T}$ by employing cross-entropy loss with the Total Variance distance: $TV(y^{D}_{i},y^{T}_{i})$, where $y^{D}_{i}$ and $y^{T}_{i}$ represent the vocabulary distribution of $M_{D}$ and $M_{T}$ respectively at the $i^{th}$ token position. We evaluate the quality of the classifier by measuring its F1 score on the validation set. The F1 results obtained on the datasets are relatively high; for instance, the F1 score on the MBPP dataset is 95% while the F1 score of the optimal static SL is 85%. For further details, see App. E and App. D.
Baselines
We compare the LLM inference latency of DISCO to both static SL (Static SL-5) and dynamic heuristic SL setups (dynHeur SL; see Fn. 2). We also consider the optimal static SL baseline tuned on our validation sets (Static SL-opt). Finally, we compare DISCO to our oracle (Sec. 2) that represents the lower bound on latency.
Results
Dataset Method Latency(ms) Speedup
MBPP Target 23.21 1.00x
dynHeur SL 20.07 1.16x
Static SL-5 15.88 1.46x
Static SL-opt 14.16 1.64x
DISCO (ours) 12.58 1.84x
Oracle 10.18 2.28x
HumanEval* Target 23.46 1.00x
dynHeur SL 22.57 1.04x
Static SL-5 14.43 1.63x
Static SL-opt 14.42 1.63x
DISCO (ours) 12.78 1.84x
Oracle 10.59 2.22x
CNN-DM Target 38.29 1.00x
dynHeur SL 21.18 1.81x
Static SL-5 19.74 1.94x
Static SL-opt 20.66 1.85x
DISCO (ours) 17.85 2.15x
Oracle 15.41 2.48x
Alpaca Target 47.67 1.00x
dynHeur SL 31.83 1.50x
Static SL-5 23.79 2.00x
Static SL-opt 23.65 2.02x
DISCO (ours) 22.49 2.12x
Oracle 20.04 2.38x
Table 1: Average latency results on different benchmarks. HumanEval results (marked with *) use a classifier trained on MBPP (transfer learning).
Table 1 presents our results. Employing an SL classifier consistently outperforms all other baselines across all benchmarks. Average latency improvements of DISCO over the optimal static SL and the dynamic heuristic baselines are 10.3% and 31.4% respectively (see App. F), while preserving the same output as the target model. Importantly, our improvement does not come only from our training data: the optimal static SL (as fit by that data) is still underperformed by DISCO. Finally, we note that classifiers transfer well across tasks: when trained on MBPP, DISCO is still able to outperform all baselines on HumanEval.
5 Related Work
There has been a growing body of work in the emerging field of speculative decoding for LLM acceleration. Pioneering studies by Leviathan et al. (2023) and Chen et al. (2023) introduced a novel rejection sampling scheme which preserves the distribution of the target model, effectively guaranteeing that speculative decoding setups will maintain the quality of the target model. Subsequent work Miao et al. (2024) elevated the average number of accepted tokens by using several draft models. More recent works Cai et al. (2024); Li et al. (2024); Bhendawade et al. (2024) eliminated the need for a separate draft model by incorporating specialized draft layers on top of the transformer decoder to forecast a sequence of draft tokens. All of the above works employed a static SL.
A very recent concurrent work by S et al. (2024) enhanced the draft model’s accuracy by granting it access to the target model’s representations. In addition, it employed a classifier to determine whether to halt or continue the speculation process. In this sense, this work is the closest to ours; however, our work delves into the impact of the SL on the efficiency of speculative decoding, encompassing comparisons between static and dynamic SL approaches, as well as the upper bound of improvement represented by the oracle SL.
6 Conclusion
We introduced DISCO, a dynamic speculation length optimization method. The method uses a classifier that determines whether the draft model should continue to generate the next token or pause and transition to the target model for validation. We evaluated DISCO’s effectiveness using four benchmarks and demonstrated average speedup gains of 10.3% and 31.4% relatively to the optimal static SL and dynamic heuristic baselines. Our results highlight the potential of further reducing inference cost by using simple, efficient techniques.
7 Limitations
While our study shows improved effectiveness of speculative decoding dynamic SL adaptation, several limitations should be acknowledged.
First, although we conducted experiments on four different datasets, our evaluation was limited to only two pairs of models, namely Vicuna and Starcoder model families. Extending the analysis to a broader range of model configurations could provide a more comprehensive understanding of the generalizability of our findings.
Second, we note that the effectiveness of dynamic speculation length (SL) approaches is contingent upon the standard deviation of the SL as measured by the oracle. When the standard deviation is low, there may be less room for improvement using dynamic SL strategies.
8 Potential Risks
As this work is dedicated to evaluating existing models on publicly available datasets, we are not aware of any potential negative impact.
Appendix A Oracle SL Analysis
Table 2 shows the oracle SL expectancy and standard deviation of the oracle measured on different datasets and models and Figure 5 shows the probability distribution of the oracle SL for the different datasets. We observe a high variance of SL values.
Datasets Target Model Draft Model Oracle SL MBPP Starcoder-15B Starcoder-168M 18.0 ($\pm$ 46.2) HumanEval Starcoder-15B Starcoder-168M 14.7 ($\pm$ 42.6) CNN-DM Vicuna-13B Vicuna-68M 03.2 ($\pm$04.0) Alpaca Vicuna-13B Vicuna-68M 02.4 ($\pm$02.1)
Table 2: Average (and STD) of the oracle SL as determined by the oracle per dataset and target/draft models.
Figure 5: Oracle SL probability histogram on the different datasets. We observe a high variance of SL values.
Fig. 6 shows bar charts of the average oracle SL over the normalized index of the Speculation Iteration. We calculate the bars as follows. For each prompt of a dataset, we have its corresponding sequence of oracle SLs. The length of the sequence is equal to the number of Speculation Iterations. For example, consider a prompt with oracle SLs $\langle 7,3,13,21,8\rangle$. Its normalized index is $\langle 0,0.25,0.5,0.75,1\rangle$. The bars represent the average oracle SL of buckets of size $0.0001$.
(a) MBPP
(b) HumanEval
(c) CNN-DM
(d) Alpaca
Figure 6: Bar chart of the average oracle SL. Note that the iteration index seems to have low predictive power for the oracle SL.
Appendix B Datasets and Prompts Details
We use standard datasets from Hugging Face and standard prompts from the state-of-of-the-art. Tab. 3 summarizes the composition of the datasets. We provide more details per dataset in the next sections.
Datasets Train Validation Test
MBPP 374 80 80
HumanEval - - 80
CNN-DM 500 80 80
Alpaca 500 80 80
Table 3: Number of samples per dataset and split.
B.1 MBPP
For MBPP, we use the ‘train’, ‘validation’ and ‘test’ splits of the ‘full’ subset. The whole ‘train’ split is used for training, while 80 randomly selected samples of the ‘validation’ and ‘test’ splits are respectively used for validation and test. MBPP is distributed under the cc-by-4.0 License.
Concerning the prompt, we followed Ben Allal et al. (2022); Fried et al. (2023) and included the description of the programming task and a single test to verify solution, in order to help the model catch the signature of the function (see Fig. 7).
"""{text}
{test_list[0]}
Figure 7: MBPP Prompt
B.2 HumanEval
HumanEval dataset contains a single subset with a single split (‘test’ split). We use 80 randomly selected samples of that split for test. Note that since we evaluate transfer learning from MBPP, we don’t need HumanEval training and validation sets. HumanEval is distributed under the MIT License.
Prompt contains only prompt field from the dataset.
B.3 CNN-DM
For CNN-DM, we use the ‘train’, ‘validation’ and ‘test’ splits of the ‘2.0.0’ subset. 500 randomly selected samples of the ‘train’ split is used for training, while 80 randomly selected samples of the ‘validation’ and ‘test’ splits are respectively used for validation and test. CNN-DM is distributed under the Apache License 2.0.
We included the article field in the prompt as in Fig. 8.
"""Summarize:
{article}
Summary:
Figure 8: CNN-DM Prompt
B.4 Alpaca
Alpaca dataset contains a single split (‘train’ split). As for CNN-DM, 500 randomly selected samples of the ‘train’ split is used for training, while 80 randomly selected samples of the ‘validation’ and ‘test’ splits are respectively used for validation and test. Alpaca is distributed under the cc-by-nc-4.0 License.
We follow Taori et al. (2023) to define the prompts. For samples with a non-empty input field, we use the prompt as in Fig. 9 while for samples with empty input field, we use the prompt as in Fig. 10.
 """Below is an instruction that describes a
 task, paired with an input that provides
 further context. Write a response that
 appropriately completes the request.
### Instruction:
{instruction}
### Input:
{input}
### Response:
Figure 9: Alpaca prompt for samples with a non-empty input field.
"""Below is an instruction that describes a
task. Write a response that appropriately
completes the request.
### Instruction:
{instruction}
### Response:
Figure 10: Alpaca prompt for samples with empty input field.
Appendix C Models
For all models, we retrieve model weights from Hugging Face. For clarity and reproducibility, we provide the URLs for each model used:
Appendix D Classifier
D.1 Classifier Training Data
In order to train the model we created training data using the following method: the target model followed the ‘standard’ autoregressive approach to generate tokens based on the input prompt, whereas the draft model iteratively generated a single token per iteration. In each iteration, the draft model generates the next token based on the concatenation of the input prompt and the tokens subsequently generated by the target model. Draft tokens resembling the target token at the same position were labeled ‘accepted’, while others were labeled ‘rejected’. Corresponding features were extracted for both draft and target tokens, encompassing the top-k probabilities of the vocabulary distribution, the entropy associated with these probabilities, and the token position value counted from the beginning of the generation process.
D.2 Classifier F1 Results
We evaluate the quality of the classifier by measuring its F1 score on the validation set and report in Table 4 F1 scores for both static SL and DISCO. F1 scores of DISCO always outperforms F1 scores of static SL.
Datasets F1 static-SL F1 DISCO
MBPP 85 95
CNN-DM 76 88
Alpaca 68 77
Table 4: Classifier F1 scores for both static SL and DISCO.
Appendix E Additional Implementation Details
Our implementation is based on the Transformers library of HuggingFace, distributed under the Apache License 2.0, and PyTorch Deep Learning library, distributed under the BSD License (BSD-3). Our code will be available upon publication under the Apache License 2.0.
For every dataset, DISCO classifier is trained on the train set; threshold $\tau$ and $SL_{max}$ hyper-parameters are fine-tuned on the validation set optimizing the latency. Optimal static SL is estimated on the validation set. Latency results are reported on the test set.
All our experiments are run on a single A100 80GB GPU.
Appendix F Additional Results
Table 5 shows the percentage of improvement in latency of DISCO over Static SL-opt and dynHeur SL baselines. The numbers are calculated based on the latency results shown in Table 1
Dataset Static SL-opt dynHeur SL
MBPP 11.2 37.3
HumanEval 11.4 43.4
CNN-DM 13.6 15.7
Alpaca 4.9 29.3
Average 10.3 31.4
Table 5: The latency improvement (percentage) of DISCO over Static SL-opt and dynHeur SL across four datasets.

### v4

@@
-Accelerating Speculative Decoding using Dynamic Speculation Length
+Dynamic Speculation Lookahead Accelerates Speculative Decoding of Large Language Models
 Abstract
-Speculative decoding is a promising method for reducing the inference latency of large language models.
-The effectiveness of the method depends on the speculation length (SL)—the number of tokens generated by the draft model at each iteration.
-The vast majority of speculative decoding approaches use the same SL for all iterations.
-In this work, we show that this practice is suboptimal.
-We introduce DISCO, a DynamIc SpeCulation length Optimization method that uses a classifier to dynamically adjust the SL at each iteration, while provably preserving the decoding quality.
-Experiments with four benchmarks demonstrate average speedup gains of 10.3% relative to our best baselines.
-[^1] Code will be available upon publication.
-Accelerating Speculative Decoding using Dynamic Speculation Length
+Speculative decoding is commonly used for reducing the inference latency of large language models.
+Its effectiveness depends highly on the speculation lookahead (SL)—the number of tokens generated by the draft model at each iteration.
+In this work we show that the common practice of using the same SL for all iterations (static SL) is suboptimal.
+We introduce DISCO (DynamIc SpeCulation lookahead Optimization), a novel method for dynamically selecting the SL.
+Our experiments with four datasets show that DISCO reaches an average speedup of 10% compared to the best static SL baseline, while generating the exact same text.
+[^1] We will publicly release our code upon publication.
+Dynamic Speculation Lookahead Accelerates
+Speculative Decoding of Large Language Models
 Jonathan Mamou  Oren Pereg  Daniel Korat Moshe BerchanskyNadav TimorMoshe WasserblatRoy Schwartz Intel Labs, Israel Weizmann Institute of Science, Israel School of Computer Science & Engineering, Hebrew University of Jerusalem Correspondence: jonathan.mamou@intel.com
 1 Introduction
-Large language models (LLMs) generate tokens in an autoregressive manner, incurring slow and costly generation.
-Speculative decoding methods Leviathan et al.
+Large language models (LLMs) generate tokens autoregressively (Radford et al., 2019; Brown et al., 2020), which often leads to slow generation.
+Speculative decoding algorithms Leviathan et al.
 (2023); Miao et al.
 (2024); Chen et al.
-(2023) mitigate this issue by splitting the inference of a given target model into two steps.
-In the first step, a faster draft model generates tokens autoregressively.
-In the second step, the target model validates all the draft tokens simultaneously.
-The validation procedure accepts or rejects the draft tokens (Fig.
-1).
-This approach accelerates the generation process by producing multiple tokens in a single forward pass of the target model, while maintaining the decoding quality of the target model.
-The effectiveness of speculative decoding in expediting the token generation process depends on the speculation length (SL) parameter, which pertains to the number of tokens generated by the draft model before each validation step (Chen et al., 2023).
-The choice of SL has a large effect on latency: lower static (constant) SL values may incur a lower number of accepted draft tokens per speculative iteration, which leads to more target model validations than necessary.
-In contrast, higher static SL may result in a large number of ignored draft generations in case of an early draft-target mismatch (Fig.
-1).
-Figure 1: An illustration of a single speculative decoding iteration with static speculation length ($SL$) $=5$.
-A draft model generates $SL$ tokens autoregressively.
-The target model validates them all in parallel, and accepts the first three tokens.
-In this case, $SL$ is too high (as the last two draft tokens are never used).
-$SL$ can also be too low (e.g., $SL=2$ here), which would lead to too many calls to the target model.
-The vast majority of speculative decoding approaches Leviathan et al.
-(2023); Miao et al.
-(2024); Chen et al.
-(2023); Cai et al.
-(2024); Li et al.
-(2024) employ a static SL, i.e., a constant number of draft tokens are generated in each speculative iteration.
-In this work, we show that this approach is suboptimal.
-We start by introducing a dynamic oracle SL, which sets the SL to be the exact number of draft generations until the first mismatch in each speculative iteration.
-We observe a high variance in the optimal SL across the speculative iterations.
-This finding demonstrates the potential of dynamically adapting the speculation length.
-We then propose DISCO, a dynamic speculation length optimization method.
-DISCO uses draft model features to train a lightweight classifier that predicts the compatibility between a generated draft token and the target token.
-At inference time, the classifier determines whether the draft model should continue generating the next token, or halt and switch to the target model for validation.
-We evaluate DISCO’s effectiveness in code generation, text summarization and instruction tasks.
-Our results demonstrate average speedup gains of 10.3% and 31.4% relatively to the best static and dynamic SL baselines, respectively, while maintaining the same generated text.
-Our classifier also transfers well across tasks from the same category: training it on one task and using it on another leads to similar gains over our baselines.
-Figure 2: The improvement factor (IF) as a function of the speculation length ($\gamma$) for three different acceptance rate ($\alpha$) values assuming fixed cost coefficient ($c=0.05$).
-The IF value largely depends on both $\alpha$ and $\gamma$.
+(2023) reduces the inference latency by splitting the inference of a given model into two steps.
+First, a fast draft model generates tokens autoregressively.
+Then a more accurate (target) model validates all generated draft tokens simultaneously.
+See Fig.
+1 for an example.
+The effectiveness of speculative decoding depends on the speculation lookahead (SL)—the number of tokens generated by the draft model at each iteration.
+An SL too small leads to too many target forwards steps; SL values too large add redundant draft forward passes.
+Yet, existing speculative decoding approaches use a static SL—an SL that remains constant across all iterations (Leviathan et al., 2023; Miao et al., 2024; Chen et al., 2023; Cai et al., 2024; Li et al., 2024).
+Figure 1: An illustration of a single speculative decoding iteration with Speculation Lookahead (SL) = 5.
+Given a prompt $t_{0}$, a draft model autoregressively generates 5 tokens $t_{1},\ldots,t_{5}$.
+The target model validates them all in parallel and accepts only $t_{1},t_{2},t_{3}$.
+As $t_{4}$ and $t_{5}$ are rejected, the SL is suboptimal (too large).
+Figure 2: Oracle and static SL values for different speculative iterations on one MBPP example.
+For static SL, we run 38 target forward passes and 192 draft forward passes, while for oracle SL, we only run 27 target forward passes and 129 draft forward passes.
+We observe a high variance of oracle SL values.
+This work starts by defining an oracle SL—a method that assigns each iteration its optimal SL.
+We observe that the optimal SL shows a high variance across iterations (Fig.
+2).
+We then use this oracle to estimate an upper bound of the expected speedup (compared to using static SLs), showing a potential gain of up to 39% speedup.
+We then propose DISCO, a novel method for selecting the SL before each iteration.
+DISCO estimates the likelihood of the next draft token being accepted by the target model, and halts the draft model if this likelihood is too small.
+We evaluate DISCO across various tasks: code generation, text summarization, and instruction following.
+Our results show an average speedup of 10% compared to optimal static SL and 31% compared to a previously known heuristic for controlling SLs Gante (2023), all without modifying the output text (Appendix F).
+DISCO also transfers well across tasks from the same category: training it on one task and using it on another leads to similar speedups.
 2 Background: Speculative Decoding
-Speculative decoding Leviathan et al.
-(2023); Miao et al.
-(2024); Chen et al.
-(2023) expedites LLM generation while ensuring no accuracy loss by dividing it into two stages.
+Speculative decoding expedites LLM generation while ensuring no accuracy loss by dividing it into two stages.
 In the first stage, a fast but less accurate draft model $M_{D}$ autoregressively generates a sequence of tokens.
@@
 This process allows the model to potentially produce multiple tokens per target forward pass.
-Speculation length (SL)
+Speculation lookahead (SL)
 The effectiveness of speculative decoding in accelerating the token generation process relies heavily on the SL parameter, which determines how many tokens are generated by the draft model before each validation step.
 The effect of the SL on the overall speedup is subject to a tradeoff; higher SLs potentially reduce the number of target model validations, but also increase the number of redundant draft generations (Fig.
-1).
+1), and vice-versa.
 The majority of speculative decoding approaches Leviathan et al.
@@
 (2024); Li et al.
-(2024) use a static SL—the same number of draft tokens are generated per speculative iteration.222 Some practical work apply heuristics for dynamic SL adaptation, increasing/decreasing SL when all/some draft tokens are accepted/rejected in the previous iteration (https://huggingface.co/blog/assisted-generation).
+(2024) use a static SL—the same number of draft tokens are generated per speculative iteration.
+[^2] A notable exception is Gante (2023), who applies a heuristic for dynamic SL adaptation by modifying the SL based on the acceptance rate of previous iterations.
 Chen et al.
 (2023) explored various static SLs, across different target-draft model pairs and tasks, and empirically showed that as the SL rises, the overall speedup increases until reaching a certain threshold, beyond which it either levels off or even regresses.
-To study the effect of different static SL values, Leviathan et al. (2023) defined the improvement factor (IF) as the expected token generation speedup. Formally:
+To study the effect of the SL, Leviathan et al.
+(2023) defined the improvement factor (IF) as the expected token generation speedup:
 $IF={\frac{1-\alpha^{\gamma+1}}{(1-\alpha)({\gamma}c+1)}}$ (1)
@@
 [^3] IF computation assumes enough compute resources for increased concurrency as $\gamma$ rises.
-Figure 2 shows the IF as a function of $\gamma$ for three different values of $\alpha$ and a fixed $c$ value.
 While both $\alpha$ and $c$ are important to the selection of the target-draft model pair, finding the optimal $\gamma$ is fundamental to the effectiveness of the system.
-3 Dynamic Speculation Length
-The IF function is based on the simplifying assumption that the probability of accepting draft tokens by the target model is i.i.d.
-Nevertheless, in practical scenarios, different tokens may have varying levels of predictability, rendering the simplifying i.i.d.
-assumption unsound.
-Therefore, relying on a static SL might not be optimal, as certain speculative iterations may benefit from generating more tokens than others.
+3 Dynamic Speculation Lookahead
+The IF function (Eq.
+1) is based on the simplifying assumption that the probability of accepting draft tokens by the target model is i.i.d.
+Nevertheless, in practical scenarios, different tokens may have varying levels of predictability, which challenges this i.i.d.
+assumption, and suggests that using a static SL might be suboptimal.
 Below we consider an oracle experiment, which applies the optimal dynamic $\gamma$ value at each iteration.
 We then propose a method for dynamically setting $\gamma$, showing that it strongly outperforms a static selection method for any choice of a static SL.
-Assessing the optimal dynamic $\gamma$ value
-We start by employing an oracle for detecting the ‘ground truth’ value of $\gamma$ for each speculative iteration.
+Finding the optimal SL per iteration
+We start by employing an oracle for detecting the optimal value of SL ($\gamma$) for each speculative iteration.
 The oracle uses the draft model to autoregressively generate tokens until a mismatch occurs between the predicted tokens of the draft and target models.
@@
 Figure 2 shows the oracle SL values across the speculative iterations for one MBPP example.
-Comparing with static SL, we observe a lower number of oracle draft and target forward passes.
-Figure 4 shows the average oracle SL over the speculative iterations for Alpaca dataset.
-We observe on both figures a high variance of oracle SL values, implying that setting a predefined static SL is likely to be suboptimal.
+Compared to the static SL, we observe a lower number of both draft and target forward passes.
+Figure 3 shows the average oracle SL over the speculative iterations for the Alpaca dataset Taori et al.
+(2023).
+Both figures show a high variance of oracle SL values, implying that astatic SL is likely to be suboptimal.
 See App.
 A for further analysis.
-Figure 3: Oracle and static SL values for different speculative iterations on one MBPP example.
-For static SL, we run 38 target forward passes and 192 draft forward passes, while for oracle SL, we only run 27 target forward passes and 129 draft forward passes.
-We observe a high variance of oracle SL values.
-Figure 4: Bar chart of the average oracle SL over the normalized index of the speculative iterations for the Alpaca dataset. We observe a high variance of oracle SL values.
-DynamIc SpeCulation length Optimization
+Figure 3: The average oracle SL over the normalized index of the speculative iterations for the Alpaca dataset. We observe a high variance of oracle SL values.
+DynamIc SpeCulation lookahead Optimization
 We introduce DISCO, a simple method for dynamically setting the SL value at each iteration.
-To estimate the correct SL value at each step, we employ a simple SL classifier.
-Our classifier decides, at each draft generation step, whether the draft model $M_{D}$ should proceed to generate the next token or switch to the target model $M_{T}$ for verification.
+To estimate the correct SL value at each step, we employ a simple classifier as follows.
+Immediately after generating any draft token, our classifier decides whether the draft model $M_{D}$ should proceed and generate the next token or switch to the target model $M_{T}$ for verification.
 The classifier takes as inputs the probability vector of the draft model ($y^{D}_{i}$) and the token position ($i$), generating a confidence score ($C_{i}$) used for the decision-making as follows:
 $C_{i}=FFN(Concat(Top_{k}(y^{D}_{i}),Ent(y^{D}_{i}),i))$
-where $Top_{k}()$ selects the top $k$ values [^4] We use $k=10$ in all experiments. and $Ent()$ is the entropy function.
+where $Top_{k}()$ selects the top $k$ values and $Ent()$ is the entropy function. [^4] We use $k=10$ in all experiments.
 […]
@@
 Datasets
-We evaluate our method on four datasets spanning three tasks: code generation, using MBPP Austin et al.
+We evaluate our method on four datasets spanning three tasks: code generation using MBPP Austin et al.
 (2021) and HumanEval Chen et al.
 (2021); text summarization using CNN-DailyMail Nallapati et al.
-(2016); and instruction-following using Alpaca Taori et al.
-(2023).
-We use the training set for training the SL classifier and the validation set for setting threshold $\tau$ and $SL_{max}$ hyperparameters.
+(2016); and instruction-following using Alpaca.
+We use the training sets for training the SL classifier and the validation sets for setting the threshold $\tau$ and the $\text{SL}_{\max}$ hyperparameter.
 For HumanEval, which has no training and validation sets, we evaluate transfer learning from MBPP.
@@
 Models
-To evaluate speculative decoding on code generation tasks, we use the Starcoder model family Li et al.
+For code generation tasks, we use the Starcoder model family Li et al.
 (2023)—15B for target and 168M for draft.
-For our text generation tasks, we use Vicuna models—13B as target Chiang et al.
+For the other tasks, we use Vicuna models—13B as target Chiang et al.
 (2023) and 68M as draft Yang et al.
 (2024).
-Throughout the experiments we use the rejection sampling scheme Leviathan et al.
-(2023) with the temperature set to zero.
-Exact model names and versions are provided in App.
-C.
+See App.
+C for more details.
 SL classifier training
-For minimal overhead, we use a shallow 2-layer FFN classifier.
-We train the SL classifier to predict the conformity between $M_{D}$ and $M_{T}$ by employing cross-entropy loss with the Total Variance distance: $TV(y^{D}_{i},y^{T}_{i})$, where $y^{D}_{i}$ and $y^{T}_{i}$ represent the vocabulary distribution of $M_{D}$ and $M_{T}$ respectively at the $i^{th}$ token position.
+To train the classifier we extract features from the training sets of our datasets MBPP, CNN-DM, and Alpaca.
+For minimal overhead, we use a shallow 2-layer FFN classifier and train it based on the extracted features to predict the agreement between the draft and target models.
+The training employs cross-entropy loss with Total Variance as distance measure: $TV(y^{D}_{i},y^{T}_{i})$, where $y^{D}_{i}$ and $y^{T}_{i}$ represent the vocabulary distribution of $M_{D}$ and $M_{T}$ respectively at the $i^{th}$ token position.
 We evaluate the quality of the classifier by measuring its F1 score on the validation set.
-The F1 results obtained on the datasets are relatively high; for instance, the F1 score on the MBPP dataset is 95% while the F1 score of the optimal static SL is 85%.
-For further details, see App.
-E and App.
-D.
+The F1 results obtained on the datasets are relatively high; for instance, 95% on MBPP, compared to 85% using the optimal static SL.
+See Apps.
+D and E for more details.
 Baselines
-We compare the LLM inference latency of DISCO to both static SL (Static SL-5) and dynamic heuristic SL setups (dynHeur SL; see Fn.
-2).
+We compare the LLM inference latency of DISCO to both static SL (static SL-5) and dynamic heuristic SL setups (dynHeur SL; Gante, 2023).
 We also consider the optimal static SL baseline tuned on our validation sets (static SL-opt).
-Finally, we compare DISCO to our oracle (Sec.
-2) that represents the lower bound on latency.
+Finally, we also report results for our oracle (Sec.
+2), which represents the lower bound on latency.
 Results
-Dataset Method Latency(ms) Speedup
+Benchmark Method Latency(ms) Speedup
 MBPP Target 23.21 1.00x
@@
 oracle 10.18 2.28x
-HumanEval* Target 23.46 1.00x
+HumanEval (transfer learning) Target 23.46 1.00x
 dynHeur SL 22.57 1.04x
@@
 oracle 20.04 2.38x
-Table 1: Average latency results on different benchmarks. HumanEval results (marked with *) use a classifier trained on MBPP (transfer learning).
-Table 1 presents our results.
+Table 1: Average latency results on different benchmarks. HumanEval results use a classifier trained on MBPP (transfer learning). All results are provided with greedy decoding (temperature=0).
+Table 1 presents our results using the rejection sampling scheme with greedy decoding (temperature=0) since baselines get higher speedup Leviathan et al.
+(2023).
 Employing an SL classifier consistently outperforms all other baselines across all benchmarks.
-Average latency improvements of DISCO over the optimal static SL and the dynamic heuristic baselines are 10.3% and 31.4% respectively (see App.
-F), while preserving the same output as the target model.
+Average latency improvements of DISCO over the optimal static SL and the dynamic heuristic baselines are 10.3% and 31.4% respectively, while preserving the same output as the target model.
 Importantly, our improvement does not come only from our training data: the optimal static SL (as fit by that data) is still underperformed by DISCO.
-Finally, we note that classifiers transfer well across tasks: when trained on MBPP, DISCO is still able to outperform all baselines on HumanEval.
+Finally, DISCO transfers well across tasks: when trained on MBPP, it is still outperforms all baselines on HumanEval.
+Analysis
+Since DISCO is sampling temperature agnostic, we present in Table 2 latency results for CNN-DM with non-zero temperature; we observe that DISCO speedup also improves in that case.
+In addition, we report latency results using a simple rule-based approach: we use the perplexity to measure the confidence in the sequence of tokens predicted by the draft, when lower perplexity indicates higher confidence.
+Optimal perplexity threshold is tuned on our validation sets (ppl SL-opt).
+We observe that it yields a lower speedup.
+Benchmark Method Latency(ms) Speedup
+CNN-DM Target 36.32 1.00x
+static SL-5 21.88 1.66x
+DISCO (ours) 19.96 1.82x
+ppl SL-opt 26.79 1.43x
+Table 2: Additional average latency results for CNN-DM: temperature=1; perplexity-based baseline.
 5 Related Work
-There has been a growing body of work in the emerging field of speculative decoding for LLM acceleration.
-Pioneering studies by Leviathan et al.
-(2023) and Chen et al.
-(2023) introduced a novel rejection sampling scheme which preserves the distribution of the target model, effectively guaranteeing that speculative decoding setups will maintain the quality of the target model.
+Pioneering studies on speculative decoding Leviathan et al.
+(2023); Chen et al.
+(2023) introduced a rejection sampling scheme that preserves the distribution of the target model, guaranteeing that speculative decoding maintains the quality of the target model.
 Subsequent work Miao et al.
 (2024) elevated the average number of accepted tokens by using several draft models.
-More recent works Cai et al.
-(2024); Li et al.
+Most recently, Timor et al.
+(2024) introduced DSI, a distributed variation of speculative decoding that is provably faster than non-distributed methods and does not require additional training or architectural changes.
+DSI orchestrates multiple servers, and the maximum number of servers is inversely proportional to the SL.
+To eliminate the need for a separate draft model, Li et al.
+(2024); Cai et al.
 (2024); Bhendawade et al.
-(2024) eliminated the need for a separate draft model by incorporating specialized draft layers on top of the transformer decoder to forecast a sequence of draft tokens.
-All of the above works employed a static SL.
+(2024); Yang et al.
+(2024) train additional, specialized draft layers on top of the transformer decoder.
+DISCO transfers well within domains and does not require classifier training per dataset whereas these methods necessitate training per dataset.
+At inference time, they employ a static SL; we believe that DISCO can be beneficially applied to these approaches, we leave this research for future work.
+Zhang et al.
+(2023) proposed draft-exiting with an adaptive threshold for self-speculative decoding using a rule-based approach that compares a confidence to a predetermined threshold.
+This method seems suitable to approaches where the draft is a subset of the target model, whereas our approach is more generic.
 A very recent concurrent work by S et al.
@@
 In addition, it employed a classifier to determine whether to halt or continue the speculation process.
-In this sense, this work is the closest to ours; however, our work delves into the impact of the SL on the efficiency of speculative decoding, encompassing comparisons between static and dynamic SL approaches, as well as the upper bound of improvement represented by the oracle SL.
+Our work delves into the impact of the SL on the efficiency of speculative decoding, encompassing comparisons between static and dynamic SL approaches, as well as the upper bound of improvement represented by the oracle SL.
 6 Conclusion
-We introduced DISCO, a dynamic speculation length optimization method.
+We have shown that using the same speculation lookahead parameter across speculative decoding iterations is suboptimal.
+We introduced DISCO, a dynamic speculation lookahead optimization method.
 The method uses a classifier that determines whether the draft model should continue to generate the next token or pause and transition to the target model for validation.
@@
 Extending the analysis to a broader range of model configurations could provide a more comprehensive understanding of the generalizability of our findings.
-Second, we note that the effectiveness of dynamic speculation length (SL) approaches is contingent upon the standard deviation of the SL as measured by the oracle.
+Second, we note that the effectiveness of dynamic speculation lookahead (SL) approaches is contingent upon the standard deviation of the SL as measured by the oracle.
 When the standard deviation is low, there may be less room for improvement using dynamic SL strategies.
@@
 Figure 4: Oracle SL probability histogram on the different datasets. We observe a high variance of SL values.
-Fig.
-6 shows bar charts of the average oracle SL over the normalized index of the Speculation Iteration.
+We hypothesized that later tokens are more predictive but eventually found only a relatively weak correlation, as Fig.
+2 and Fig.
+5 show.
+The figures are bar charts of the average oracle SL over the normalized index of the Speculation Iteration.
 We calculate the bars as follows.
@@
 (2023) and included the description of the programming task and a single test to verify solution, in order to help the model catch the signature of the function (see Fig.
-7).
+6).
 """{text}
@@
 Appendix D Classifier
-D.1 Classifier Training Data
-In order to train the model we created training data using the following method: the target model followed the ‘standard’ autoregressive approach to generate tokens based on the input prompt, whereas the draft model iteratively generated a single token per iteration.
-In each iteration, the draft model generates the next token based on the concatenation of the input prompt and the tokens subsequently generated by the target model.
-Draft tokens resembling the target token at the same position were labeled ‘accepted’, while others were labeled ‘rejected’.
-Corresponding features were extracted for both draft and target tokens, encompassing the top-k probabilities of the vocabulary distribution, the entropy associated with these probabilities, and the token position value counted from the beginning of the generation process.
+D.1 Feature Extraction from Training Data for Classifier Training
+To train the model for each dataset, MBPP, CNN-DM, and Alpaca, we used the corresponding training set of each dataset.
+For each token in each training set, we extracted a boolean label (accepted/rejected) and a list of features in the following manner: We ran the target model using the standard autoregressive approach for generating tokens based on the input prompt.
+In contrast, the draft model iteratively generated only a single token per iteration.
+During each iteration, the draft model generated this token based on the concatenation of the input prompt and the tokens subsequently generated by the target model.
+Draft tokens that resembled the target token at the same position were labeled "accepted," while others were labeled "rejected." Corresponding features were extracted for both draft and target tokens, encompassing the top-k probabilities of the vocabulary distribution, the entropy associated with these probabilities, and the token position value counted from the beginning of the generation process.
 D.2 Classifier F1 Results
@@
 Our code will be available upon publication under the Apache License 2.0.
-For every dataset, DISCO classifier is trained on the train set; threshold $\tau$ and $SL_{max}$ hyper-parameters are fine-tuned on the validation set optimizing the latency.
+For every dataset, DISCO classifier is trained on the train set; threshold $\tau$ and $\text{SL}_{\max}$ hyper-parameters are fine-tuned on the validation set optimizing the latency.
 Optimal static SL is estimated on the validation set.

### v5

@@
 Our experiments with four datasets show that DISCO reaches an average speedup of 10% compared to the best static SL baseline, while generating the exact same text.
-[^1] We will publicly release our code upon publication.
-Dynamic Speculation Lookahead Accelerates
-Speculative Decoding of Large Language Models
-Jonathan Mamou  Oren Pereg  Daniel Korat Moshe BerchanskyNadav TimorMoshe WasserblatRoy Schwartz Intel Labs, Israel Weizmann Institute of Science, Israel School of Computer Science & Engineering, Hebrew University of Jerusalem Correspondence: jonathan.mamou@intel.com
 1 Introduction
@@
 Yet, existing speculative decoding approaches use a static SL—an SL that remains constant across all iterations (Leviathan et al., 2023; Miao et al., 2024; Chen et al., 2023; Cai et al., 2024; Li et al., 2024).
-Figure 1: An illustration of a single speculative decoding iteration with Speculation Lookahead (SL) = 5.
-Given a prompt $t_{0}$, a draft model autoregressively generates 5 tokens $t_{1},\ldots,t_{5}$.
-The target model validates them all in parallel and accepts only $t_{1},t_{2},t_{3}$.
-As $t_{4}$ and $t_{5}$ are rejected, the SL is suboptimal (too large).
-Figure 2: Oracle and static SL values for different speculative iterations on one MBPP example.
-For static SL, we run 38 target forward passes and 192 draft forward passes, while for oracle SL, we only run 27 target forward passes and 129 draft forward passes.
-We observe a high variance of oracle SL values.
 This work starts by defining an oracle SL—a method that assigns each iteration its optimal SL.
 We observe that the optimal SL shows a high variance across iterations (Fig.
-2).
+3).
 We then use this oracle to estimate an upper bound of the expected speedup (compared to using static SLs), showing a potential gain of up to 39% speedup.
@@
 We evaluate DISCO across various tasks: code generation, text summarization, and instruction following.
-Our results show an average speedup of 10% compared to optimal static SL and 31% compared to a previously known heuristic for controlling SLs Gante (2023), all without modifying the output text (Appendix F).
+Our results show an average speedup of 10% compared to optimal static SL and 31% compared to a previously known heuristic for controlling SLs Gante (2023), all without modifying the output text (App.
+F).
 DISCO also transfers well across tasks from the same category: training it on one task and using it on another leads to similar speedups.
@@
 This process allows the model to potentially produce multiple tokens per target forward pass.
+Figure 1: An illustration of a single speculative decoding iteration with Speculation Lookahead (SL) = 5.
+Given a prompt $t_{0}$, a draft model autoregressively generates 5 tokens $t_{1},\ldots,t_{5}$.
+The target model validates them all in parallel and accepts only $t_{1},t_{2},t_{3}$.
+As $t_{4}$ and $t_{5}$ are rejected, the SL is suboptimal (too large).
 Speculation lookahead (SL)
@@
 1), and vice-versa.
-The majority of speculative decoding approaches Leviathan et al.
-(2023); Miao et al.
-(2024); Chen et al.
-(2023); Cai et al.
-(2024); Li et al.
-(2024) use a static SL—the same number of draft tokens are generated per speculative iteration.
+The majority of speculative decoding approaches use a static SL—the same number of draft tokens are generated per speculative iteration.
 [^1] A notable exception is Gante (2023), who applies a heuristic for dynamic SL adaptation by modifying the SL based on the acceptance rate of previous iterations.
@@
 We then propose a method for dynamically setting $\gamma$, showing that it strongly outperforms a static selection method for any choice of a static SL.
+Figure 2: Oracle and static SL values for different speculative iterations on one MBPP example.
+For static SL, we run 38 target forward passes and 192 draft forward passes, while for oracle SL, we only run 27 target forward passes and 129 draft forward passes.
+We observe a high variance of oracle SL values.
+Figure 3: The average oracle SL over the normalized index of the speculative iterations for the Alpaca dataset. We observe a high variance of oracle SL values.
 Finding the optimal SL per iteration
@@
 A for further analysis.
-Figure 3: The average oracle SL over the normalized index of the speculative iterations for the Alpaca dataset. We observe a high variance of oracle SL values.
 DynamIc SpeCulation lookahead Optimization
@@
 The classifier takes as inputs the probability vector of the draft model ($y^{D}_{i}$) and the token position ($i$), generating a confidence score ($C_{i}$) used for the decision-making as follows:
-$C_{i}=FFN(Concat(Top_{k}(y^{D}_{i}),Ent(y^{D}_{i}),i))$
-where $Top_{k}()$ selects the top $k$ values and $Ent()$ is the entropy function. [^4] We use $k=10$ in all experiments.
+$C_{i}=FFN(Concat(Top_{k}(y^{D}_{i}),Ent(y^{D}_{i}),i))$ (2)
 […]
 4 Experiments
-Datasets
+Datasets and Models
 We evaluate our method on four datasets spanning three tasks: code generation using MBPP Austin et al.
@@
 For HumanEval, which has no training and validation sets, we evaluate transfer learning from MBPP.
-See App.
-B for more details on the datasets and prompts.
-Models
 For code generation tasks, we use the Starcoder model family Li et al.
@@
 (2024).
-See App.
-C for more details.
+See Apps.
+B and C for more details.
 SL classifier training
@@
 D and E for more details.
-Baselines
+Baselines and Results
 We compare the LLM inference latency of DISCO to both static SL (static SL-5) and dynamic heuristic SL setups (dynHeur SL; Gante, 2023).
-We also consider the optimal static SL baseline tuned on our validation sets (static SL-opt).
-Finally, we also report results for our oracle (Sec.
-2), which represents the lower bound on latency.
-Results
-Benchmark Method Latency(ms) Speedup
+Benchmark Method Latency Speedup
 MBPP Target 23.21 1.00x
@@
 oracle 20.04 2.38x
-Table 1: Average latency results on different benchmarks. HumanEval results use a classifier trained on MBPP (transfer learning). All results are provided with greedy decoding (temperature=0).
+Table 1: Average latency results (in milliseconds) on different benchmarks.
+HumanEval results use a classifier trained on MBPP (transfer learning).
+All results are provided with greedy decoding (temperature=0).
+We also consider the optimal static SL baseline tuned on our validation sets (static SL-opt).
+Finally, we also report results for our oracle (Sec.
+2), which represents the lower bound on latency.
 Table 1 presents our results using the rejection sampling scheme with greedy decoding (temperature=0) since baselines get higher speedup Leviathan et al.
@@
 Finally, DISCO transfers well across tasks: when trained on MBPP, it is still outperforms all baselines on HumanEval.
-Analysis
-Since DISCO is sampling temperature agnostic, we present in Table 2 latency results for CNN-DM with non-zero temperature; we observe that DISCO speedup also improves in that case.
-In addition, we report latency results using a simple rule-based approach: we use the perplexity to measure the confidence in the sequence of tokens predicted by the draft, when lower perplexity indicates higher confidence.
-Optimal perplexity threshold is tuned on our validation sets (ppl SL-opt).
-We observe that it yields a lower speedup.
-Benchmark Method Latency(ms) Speedup
-CNN-DM Target 36.32 1.00x
-static SL-5 21.88 1.66x
-DISCO (ours) 19.96 1.82x
-ppl SL-opt 26.79 1.43x
-Table 2: Additional average latency results for CNN-DM: temperature=1; perplexity-based baseline.
+See App.
+G for further analysis.
 5 Related Work
@@
 (2024) introduced DSI, a distributed variation of speculative decoding that is provably faster than non-distributed methods and does not require additional training or architectural changes.
-DSI orchestrates multiple servers, and the maximum number of servers is inversely proportional to the SL.
 To eliminate the need for a separate draft model, Li et al.
@@
 Our results highlight the potential of further reducing inference cost by using simple, efficient techniques.
-7 Limitations
-While our study shows improved effectiveness of speculative decoding dynamic SL adaptation, several limitations should be acknowledged.
-First, although we conducted experiments on four different datasets, our evaluation was limited to only two pairs of models, namely Vicuna and Starcoder model families.
-Extending the analysis to a broader range of model configurations could provide a more comprehensive understanding of the generalizability of our findings.
-Second, we note that the effectiveness of dynamic speculation lookahead (SL) approaches is contingent upon the standard deviation of the SL as measured by the oracle.
-When the standard deviation is low, there may be less room for improvement using dynamic SL strategies.
-8 Potential Risks
-As this work is dedicated to evaluating existing models on publicly available datasets, we are not aware of any potential negative impact.
 Appendix A Oracle SL Analysis
-Table 3 shows the oracle SL expectancy and standard deviation of the oracle measured on different datasets and models and Figure 4 shows the probability distribution of the oracle SL for the different datasets.
+Table 2 shows the oracle SL expectancy and standard deviation of the oracle measured on different datasets and models and Fig.
+4 shows the probability distribution of the oracle SL for the different datasets.
 We observe a high variance of SL values.
@@
 Concerning the prompt, we followed Ben Allal et al.
-(2022); Fried et al.
-(2023) and included the description of the programming task and a single test to verify solution, in order to help the model catch the signature of the function (see Fig.
+[2022], Fried et al.
+[2023] and included the description of the programming task and a single test to verify solution, in order to help the model catch the signature of the function (see Fig.
 6).
@@
 Alpaca is distributed under the cc-by-nc-4.0 License.
-We follow Taori et al. (2023) to define the prompts. For samples with a non-empty input field, we use the prompt as in Fig. 8 while for samples with empty input field, we use the prompt as in Fig. 9.
+We follow Taori et al. [2023] to define the prompts. For samples with a non-empty input field, we use the prompt as in Fig. 8 while for samples with empty input field, we use the prompt as in Fig. 9.
  """Below is an instruction that describes a
@@
 D.2 Classifier F1 Results
-We evaluate the quality of the classifier by measuring its F1 score on the validation set and report in Table 5 F1 scores for both static SL and DISCO.
+We evaluate the quality of the classifier by measuring its F1 score on the validation set and report in Tab.
+4 F1 scores for both static SL and DISCO.
 F1 scores of DISCO always outperforms F1 scores of static SL.
+F1 score measures the accuracy of the classifier in predicting the speculative length (SL) but does not account for how well the predictions align with the oracle’s behavior in reducing latency.
+In particular, F1 is influenced by 2 different types of error FP and FN that have a different impact on the speedup.
+FP Errors lead to unnecessary speculative execution, which wastes resources but might not drastically reduce overall speedup.
+FN Errors lead to missed opportunities for speedup, having a more severe impact on overall latency reduction.
 Datasets F1 static-SL F1 DISCO
@@
 Table 5: The latency improvement (percentage) of DISCO over static SL-opt and dynHeur SL across four datasets.
+Appendix G Further Latency Results Analysis
+Concerning SL-opt and SL-5 speedup values on CNN-DM reported in Tab.
+1, note that the optimal static SL is tuned on the validation set while the latency and speedup reported numbers are on the test set.
+This is similar to our setup, where the classifier is tuned on the validation set, and similarly reported on the test set.
+We observe that for CNN-DM, the optimal SL on the validation set is not as good as the SL-5 on the test set.
+Since DISCO is sampling temperature agnostic, we present in Tab.
+6 latency results for CNN-DM with non-zero temperature; we observe that DISCO speedup also improves in that case.
+In addition, we report latency results using a simple rule-based approach: we use the perplexity to measure the confidence in the sequence of tokens predicted by the draft, when lower perplexity indicates higher confidence.
+Optimal perplexity threshold is tuned on our validation sets (ppl SL-opt).
+We observe that it yields a lower speedup.
+Benchmark Method Latency(ms) Speedup
+CNN-DM Target 36.32 1.00x
+static SL-5 21.88 1.66x
+DISCO (ours) 19.96 1.82x
+ppl SL-opt 26.79 1.43x
+Table 6: Additional average latency results for CNN-DM: temperature=1; perplexity-based baseline.